(d^2+dd'+d'+1) =5e^x


1
Expert's answer
2021-05-17T02:49:01-0400

We have given the differential equation:


"(D^2+DD'+D'+1)z = 5e^x"


The given equation cannot be written as linear factors. Hence, it's complimentary function is taken as a trial solution


"z = \sum Ae^{hx+ky}"


Therefore we have,


"Dz = \sum Ahe^{hx+ky}"


"D'z = \sum Ake^{hx+ky}"


"DD'z = \sum Ahke^{hx+ky}"


Put all these values in the given equation "(D^2+DD'+D'+1)z = 0" , We have


"\sum Ah^2e^{hx+ky}+\sum Ahke^{hx+ky}+\sum Ake^{hx+ky}+\sum Ae^{hx+ky} = 0"


"\sum A(h^2+hk+k+1)e^{hx+ky} = 0 \implies (h^2+hk+k+1) = 0"


"k = -\dfrac{(h^2+1)}{(h+1)}"


Thus,


"CF = \sum Ae^{hx+ky}, where k = -\dfrac{(h^2+1)}{(h+1)}"


"PI = \dfrac{1}{F(D,D')}e^{ax+by}"


"PI = 5\dfrac{1}{D^2+DD'+D'+1}e^x"


"PI = 5\dfrac{1}{1^2+1.0+0+1}e^x"


"PI = \dfrac{5e^x}{2}"


Hence, the required solution is "= CF+PI"


"= \sum Ae^{hx+ky}+\dfrac{5e^x}{2}, where\hspace{2mm} k = -\dfrac{(h^2+1)}{(h+1)}"